x^2+(x+21)^2=(x+24)^2
Answer by Nate(3500)

x^2 + x^2 + 42x + 441 = x^2 + 48x + 576
x^2 - 6x - 135 = 0
x^2 - 6x = 135
(x - 3)^2 = 144
x - 3 = +- 12
x = 3 +- 12
x = 15 and x = -9